(a)   The market demand and supply equations for theme park in a city are given by P = 30 – 0.005QD and P = 10 + 0.005QS, where P is the price in dollars and QD is the quantity of theme-park tickets demanded and QS is the quantity of theme park ticket supplied.

i) Solve for the equilibrium price and quantity of theme park tickets and compute the consumer surplus and producer surplus. Support your answers with a graph of the theme park tickets market.

(ii) Explain the implications of the welfare of consumers, producers and the society when the price of theme park ticket is fixed at $15. Support your answers with a graph of the theme park tickets market.


1
Expert's answer
2022-02-13T11:50:19-0500

(i) In equilibrium Pd = Ps, so:

30 – 0.005Q = 10 + 0.005Q,

0.01Q = 20,

Q = 2,000 units,

P = 10 + 0.005×2,000 = 20.

(ii) If P = $15, which is below equilibrium, then the deficit of tickets will occur and the welfare will decrease.